Erythromycin 48 g in divided doses, is ordered for a patient with syphillis. You are told to give the medicine 4 times per day for 12 days. You have only 400 mg tablets. How many tablets will you give the patient each time? .
Answer #1

Ans. Erythromycin 48g, is divided doses

4 times per day , 48/4 = 12g per times

So, 400mg tablets is convert in gram 400/1000= 0.4gram

So 0.4multiply 3/10 =12gm each times 3 tablet